OpenStudy (anonymous):

Find h'(1) given that f(1) = -6, f '(1) = -5, g(1) = 5, and g'(1) = 6. (a) h(x) = 3f(x) - 3g(x)

OpenStudy (anonymous):

h'(x) =3f'(x)-3g'(x) h'(1) =3f'(1)-3g'(1) h'(1) =3(-6)-3(6)= -18 - 18 = -36

OpenStudy (anonymous):

correction* 3(-5)- 3(6) = -15 - 18= -33

OpenStudy (anonymous):

ooo i thnk i was over analyzing the question

OpenStudy (anonymous):

thank you